What is X if, logx^(16)=-2
I bet it's not -8 that would be too easy. My mind is spinning, help please!!!

Answer by Alan3354(69443)

What is X if, logx^(16)=-2
Is it log(x^16) ? If so,
log(x^16) = -2
16log(x) = -2
log(x) = -1/8
x = 10^(-1/8)
x = ~0.74989